OpenStudy (anonymous):

6a+5 esponent 2 ;a=4 6a+5 esponent 2 ;a=4 @Mathematics

OpenStudy (anonymous):

(6a + 5)^2 = (24 = 5)^2 =29^2

OpenStudy (anonymous):

So what's 29 squared?

OpenStudy (precal):

You mean (24+5)^2

OpenStudy (anonymous):

\[5^{2}\]

OpenStudy (anonymous):

yes sorry, its 25 + 4

OpenStudy (anonymous):

\[6a+5^2\]

OpenStudy (anonymous):

49